# ATLAS: Adaptive Self-Evolutionary Research Agent with Task-Distributed Multi-LLM Supporters

### v2

@@ A.1 Scope and Limitation of the Theoretical Abstraction
The theorem should be read as evidence that systematic reference refresh can yield sublinear dynamic regret in a drifting preference environment, whereas a fixed reference can incur linear bias under drift; the gated implementation aims to approximate systematic refresh while enforcing conservative stability constraints, whose reliability we quantify empirically (App. B.6B.7).
A.2 Preliminaries and Definitions
We first formalize the sliding-window estimator and its requisite assumptions. For brevity, write $\phi_{\tau}:=\phi(x_{\tau},y_{\tau})$. Let $\mathcal{W}_{t}$ denote the sliding window of recent history with size $W=|\mathcal{W}_{t}|=T^{\kappa}$ where $\kappa\in(0,1)$ and controllable parameter by decision-maker. We define the regularized covariance matrix $A_{t}$ as:
$A_{t}=\sum_{\tau\in\mathcal{W}_{t}}\phi_{\tau}\phi^{\top}_{\tau}+\lambda I_{d}$ (20)
Estimator
Let $p_{\tau}\in\{0,1\}$ denote the observed preference label at $\tau$. We use the regularized logistic regression estimator:
$\hat{\theta}_{t}=\arg\min_{\theta_{t}}\sum_{\tau\in\mathcal{W}_{t}}\ell\!\left(p_{\tau},\langle\theta_{t},\phi_{\tau}\rangle\right)+\frac{\lambda}{2}\|\theta_{t}\|_{2}^{2}$ (21)
where
$\ell\left(p_{\tau},\langle\theta_{t},\phi_{\tau}\rangle\right)=\log\left(1+e^{\theta_{t}^{\top}\phi_{\tau}}\right)-p_{\tau}\,\theta_{t}^{\top}\phi_{\tau}$ (22)
@@ A.2 Preliminaries and Definitions
$y^{*}_{t}(x)\in\arg\max_{y\in\mathcal{Y}}u_{t}(x,y)$
Assumption A.1 (Covariance Diversity).
@@
$\lambda_{\min}(A_{t})\geq c\cdot W$
Assumption A.2.
@@
$|\sigma(a)-\sigma(b)|\leq L_{\sigma}|a-b|$
A.3 Auxiliary Lemmas
Lemma A.3 (Oracle Switching Budget).
Under the margin condition with gap $\gamma>0$, the number of switches of the oracle-optimal action satisfies
@@ A.3 Auxiliary Lemmas
Fix $t\geq 2$ and abbreviate $y_{t}^{*}:=y_{t}^{*}(x_{t})$ and $y_{t-1}^{*}:=y_{t-1}^{*}(x_{t})$. This oracle-optimal action is defined in A.2. If $y_{t}^{*}=y_{t-1}^{*}$, there is nothing to show. Suppose $y_{t}^{*}\neq y_{t-1}^{*}$. By optimality at time $t$,
@@
equivalently $u_{t-1}(x_{t},y_{t}^{*})-u_{t-1}(x_{t},y_{t-1}^{*})\leq-\gamma$. Thus, $\gamma\leq 2\phi_{\max}\|\theta_{t}-\theta_{t-1}\|_{2}$
@@
Lemma A.4 (Local variation bound).
@@
$\sum_{\tau\in\mathcal{W}_{t}}\|\theta_{t}-\theta_{\tau}\|_{2}\leq WV_{t,W},\quad\sum^{T}_{t=1}V_{t,W}\leq WV_{T}.$
@@
Lemma A.5 (Self-normalized inequality).
Based on Assumption 1, $\|\phi_{\tau}\|_{2}\leq\phi_{max}$ and $\eta_{\tau}$ are zero-mean 1-sub-Gaussian. Then with probability at least $1-\delta$,
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{2}\leq\sqrt{\lambda+W\phi_{\max}^{2}}\;\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}.$
@@
We use self-normalized inequality (Lemma 9) derivation from (Abbasi-yadkori et al., 2011).
$\log\!\left(\frac{\det(A_{t})^{1/2}\det(\lambda I)^{-1/2}}{\delta}\right)=\frac{1}{2}\log\!\left(\frac{\det(A_{t})}{\det(\lambda I)}\right)+\log\!\left(\frac{1}{\delta}\right).$
Using the determinant bound,
$\det(A_{t})\leq\left(\lambda+\frac{W\phi_{\max}^{2}}{d}\right)^{d},\qquad\det(\lambda I)=\lambda^{d},$
hence
$\frac{1}{2}\log\!\left(\frac{\det(A_{t})}{\det(\lambda I)}\right)\leq\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right).$
By combining determinant-trace bound condition, we get follwoing inequality by self-normalized vector martingale with probability at least $1-\delta$
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{A_{t}^{-1}}\leq\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}.$
Convert to Euclidean norm:
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{2}\leq\sqrt{\lambda_{\max}(A_{t})}\;\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}$
Substitute $\lambda_{\max}(A_{t})\leq\lambda+W\phi_{\max}^{2}$:
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{2}\leq\sqrt{\lambda+W\phi_{\max}^{2}}\;\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\quad w.p\quad 1-\delta.$
Lemma A.6 (Parameter estimation error bound).
We use Assumption 1, A.1, A.2, Let $\eta_{\tau}=p_{\tau}-\sigma\bigl(\langle\theta_{\tau},\phi_{\tau}\rangle\bigr)$ be zero-mean 1-sub-Gaussian noise. Then with probability at least $1-\delta$,
$\|\hat{\theta}_{t}-\theta_{t}\|_{2}\leq\left[\frac{\phi_{\max}^{2}L_{\sigma}V_{t,W}}{m_{0}\cdot c}\;+\;\frac{\sqrt{\lambda+W\phi_{\max}^{2}}}{m_{0}cW}\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\;+\;\frac{\lambda\theta_{\max}}{m_{0}cW}\right]$
@@
Define the objective
$L_{t}(\theta_{t})=\sum_{\tau\in\mathcal{W}_{t}}\ell\left(p_{\tau},\langle\theta_{t},\phi_{\tau}\rangle\right)+\frac{\lambda}{2}\|\theta_{t}\|_{2}^{2}$
We use the chain rule on $L_{t}(\theta_{t})$ to get the gradient and Hessian. Then
$\nabla L_{t}(\theta_{t})=\sum_{\tau\in\mathcal{W}_{t}}\bigl(\sigma(\langle\theta_{t},\phi_{\tau}\rangle)-p_{\tau}\bigr)\,\phi_{\tau}+\lambda\theta_{t}.$
$\nabla^{2}L_{t}(\theta_{t})=\sum_{\tau\in\mathcal{W}_{t}}\sigma(\langle\theta_{t},\phi_{\tau}\rangle)\bigl(1-\sigma(\langle\theta_{t},\phi_{\tau}\rangle)\bigr)\,\phi_{\tau}\phi_{\tau}^{\top}+\lambda I$
@@
By Assumptions A.1, A.2, $\lambda_{min}(\nabla^{2}L_{t}(\theta_{t}))\geq m_{0}cW$. Using strong convexity
$\big\langle\nabla L_{t}(\hat{\theta}_{t})-\nabla L_{t}(\theta_{t}),\;\hat{\theta}_{t}-\theta_{t}\big\rangle\geq\lambda_{\min}\!\big(\nabla^{2}L_{t}(\theta_{t})\big)\cdot\|\hat{\theta}_{t}-\theta_{t}\|_{2}^{2}\geq m_{0}cW\,\|\hat{\theta}_{t}-\theta_{t}\|_{2}^{2}.$
Since the optimal $\hat{\theta}_{t}$ is computed by $\nabla_{\theta}L_{t}(\hat{\theta}_{t})=0$, we substitute this optimality to strong convexity inequality above. Then, we get following inequality
$m_{0}cW\,\|\hat{\theta}_{t}-\theta_{t}\|_{2}^{2}\;\leq\;\big\langle-\nabla L_{t}(\theta_{t}),\;\hat{\theta}_{t}-\theta_{t}\big\rangle\;\leq\;\|\nabla L_{t}(\theta_{t})\|_{2}\cdot\|\hat{\theta}_{t}-\theta_{t}\|_{2}.$
Therefore, the upper bound of estimate error is:
$\|\hat{\theta}_{t}-\theta_{t}\|_{2}\leq\frac{1}{m_{0}cW}\,\|\nabla L_{t}(\theta_{t})\|_{2}.$
Rewrite using drift and noise:
$\nabla L_{t}(\theta_{t})=\sum_{\tau\in\mathcal{W}_{t}}\Bigl[\sigma\big(\langle\theta_{t},\phi_{\tau}\rangle\big)-\sigma\big(\langle\theta_{\tau},\phi_{\tau}\rangle\big)\Bigr]\phi_{\tau}-\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}+\lambda\theta_{t},$
Take norms and apply triangle inequality:
$\|\nabla L_{t}(\theta_{t})\|_{2}\leq\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\Bigl[\sigma\big(\langle\theta_{t},\phi_{\tau}\rangle\big)-\sigma\big(\langle\theta_{\tau},\phi_{\tau}\rangle\big)\Bigr]\phi_{\tau}\Bigr\|_{2}+\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\phi_{\tau}\Bigr\|_{2}+\lambda\|\theta_{t}\|_{2}.$
Since $\sigma(\cdot)$ is $L_{\sigma}$-Lipschitz, the drift term satisfies
$\Bigl\|\sum_{\tau\in\mathcal{W}_{t}}\Bigl[\sigma\big(\langle\theta_{t},\phi_{\tau}\rangle\big)-\sigma\big(\langle\theta_{\tau},\phi_{\tau}\rangle\big)\Bigr]\phi_{\tau}\Bigr\|_{2}\leq\sum_{\tau\in\mathcal{W}_{t}}L_{\sigma}\,\|\theta_{t}-\theta_{\tau}\|_{2}\,\|\phi_{\max}\|_{2}^{2}\leq L_{\sigma}\phi^{2}_{max}W\;V_{t,W}$
Now, we bound the second term via the (Li et al., 2010) log-determinant argument. As we defined the $A_{t}$ matrix in (16),
$\lambda_{\max}(A_{t})\leq\lambda+\lambda_{\max}\!\left(\sum_{\tau\in\mathcal{W}_{t}}\phi_{\tau}\phi_{\tau}^{\top}\right)\leq\lambda+\sum_{\tau\in\mathcal{W}_{t}}\|\phi_{\tau}\|_{2}^{2}\leq\lambda+W\phi_{\max}^{2}.$
The noise term is bounded by Lemma A.5, and $\lambda\|\theta_{t}\|_{2}\leq\lambda\theta_{\max}$.
@@
$\|\nabla L_{t}(\theta_{t})\|_{2}\leq\left[{L_{\sigma}\phi_{\max}^{2}\,W\,V_{t,W}}\;+\;\sqrt{\lambda+W\phi_{\max}^{2}}\,\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\;+\;\lambda\theta_{\max}\right]$
Dividing by $m_{0}cW$ yields the result.
$\|\hat{\theta}_{t}-\theta_{t}\|_{2}\leq\left[\frac{L_{\sigma}\phi_{\max}^{2}\,V_{t,W}}{m_{0}\cdot c}\;+\;\frac{\sqrt{\lambda+W\phi_{\max}^{2}}}{m_{0}cW}\sqrt{2\left(\frac{d}{2}\log\!\left(1+\frac{W\phi_{\max}^{2}}{d\lambda}\right)+\log\!\left(\frac{1}{\delta}\right)\right)}\;+\;\frac{\lambda\theta_{\max}}{m_{0}cW}\right].$
Lemma A.7 (KL divergence bound).
Let $u_{t}(x_{t},y_{t})=\langle\theta_{t},\phi(x_{t},y_{t})\rangle$ and $\hat{u}_{t}(x_{t},y_{t})=\langle\hat{\theta}_{t},\phi(x_{t},y_{t})\rangle$. The KL-divergence between KL-regularized policy $\pi^{kl}_{t}$ and the agent policy $\pi_{t}$ is bounded by parameter estimate error:
@@
$D_{KL}(\pi^{kl}_{t}(\cdot|x_{t})||\pi_{t}(\cdot|x_{t}))\leq\frac{\phi_{max}^{2}\cdot\|\theta_{t}-\hat{\theta}_{t}\|_{2}^{2}}{2\beta^{2}}.$
@@
We use the policy definition, which is defined in (Rafailov et al., 2024)
$\pi^{kl}_{t}(y_{t}|x_{t})=\frac{1}{Z(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}u_{t}(x_{t},y_{t})}$
$Z(x_{t})=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta}u_{t}(x_{t},y_{t}))$
$\pi_{t}(y_{t}|x_{t})=\frac{1}{\hat{Z}(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}\hat{u}_{t}(x_{t},y_{t})}$
$\hat{Z}(x_{t})=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta}\hat{u}_{t}(x_{t},y_{t}))$
From the KL-divergence definition above, we derive the log term first.
$\log{\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}}=\log\frac{\frac{1}{Z(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}u_{t}(x_{t},y_{t})}}{\frac{1}{\hat{Z}(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta}\hat{u}_{t}(x_{t},y_{t})}}$
$\log\frac{\hat{Z}(x_{t})}{Z(x_{t})}+\frac{1}{\beta}(u_{t}(x_{t},y_{t})-\hat{u}_{t}(x_{t},y_{t}))$

### v3

@@ B.1 Scope and Limitation of the Theoretical Abstraction
The theorem should be read as evidence that systematic reference refresh can yield sublinear dynamic regret in a drifting preference environment, whereas a fixed reference can incur linear bias under drift; the gated implementation aims to approximate systematic refresh while enforcing conservative stability constraints, whose reliability we quantify empirically (App. C.6C.7).
B.2 Preliminaries and Definitions
We first formalize the sliding-window estimator and its requisite assumptions. Let $W_{t}$ denote the sliding window of recent comparison rounds with size $W=|W_{t}|=T^{\kappa}$, where $\kappa\in(0,1)$.
For each comparison round $\tau$, the learner observes a context $x_{\tau}$, two candidate actions $y_{\tau}^{(1)},y_{\tau}^{(2)}$, and a binary preference label
$p_{\tau}=\mathbf{1}\{y_{\tau}^{(1)}\succ y_{\tau}^{(2)}\}.$
We define the pairwise feature difference
$\psi_{\tau}=\phi(x_{\tau},y_{\tau}^{(1)})-\phi(x_{\tau},y_{\tau}^{(2)}).$
Under the Bradley–Terry preference model,
$\mathbb{P}(p_{\tau}=1\mid x_{\tau},y_{\tau}^{(1)},y_{\tau}^{(2)})=\sigma(\langle\theta_{\tau},\psi_{\tau}\rangle).$
The sliding-window estimator is
$\widehat{\theta}_{t}=\arg\min_{\theta}\sum_{\tau\in W_{t}}\left[\log(1+\exp(\theta^{\top}\psi_{\tau}))-p_{\tau}\theta^{\top}\psi_{\tau}\right]+\frac{\lambda}{2}\|\theta\|_{2}^{2}.$
The corresponding regularized covariance matrix is
$A_{t}=\sum_{\tau\in W_{t}}\psi_{\tau}\psi_{\tau}^{\top}+\lambda I_{d}.$
Since $\|\phi(x,y)\|_{2}\leq\phi_{\max}$, we have
$\|\psi_{\tau}\|_{2}\leq 2\phi_{\max}.$
@@ B.2 Preliminaries and Definitions
$y^{*}_{t}(x)\in\arg\max_{y\in\mathcal{Y}}u_{t}(x,y).$
Assumption B.1 (Covariance Diversity).
@@
$\lambda_{\min}\left(\sum_{\tau\in W_{t}}\psi_{\tau}\psi_{\tau}^{\top}+\lambda I_{d}\right)\geq cW.$
Assumption B.2 (Uniform logistic curvature).
@@
$|\sigma(a)-\sigma(b)|\leq L_{\sigma}|a-b|.$
Assumption B.3 (Boundedness).
We assume the feature map, parameter, and utilities are uniformly bounded,
$\|\phi(x,y)\|_{2}\leq\phi_{max},\quad\|\theta_{t}\|_{2}\leq\theta_{max},\quad|u_{t}(x,y)|\leq u_{max}.$
Assumption B.4 (Comparator support).
We assume the KL-regularized comparator has full support, uniformly lower-bounded by $\pi_{\min}^{{kl}}>0$:
$\pi_{t}^{{kl}}(y\mid x)\geq\pi_{\min}^{{kl}}$
for all relevant $t,x,y$. This assumption can be derived from a uniformly supported reference policy and bounded utilities for the exponential-tilted comparator, but we state it directly for clarity.
Assumption B.5 (Uniform contextual margin).
There exists $\gamma>0$ such that for every time $t$ and every relevant context $x$, the optimal action
$y_{t}^{*}(x)\in\arg\max_{y\in\mathcal{Y}}u_{t}(x,y)$
is unique and satisfies
$u_{t}(x,y_{t}^{*}(x))-\max_{y\neq y_{t}^{*}(x)}u_{t}(x,y)\geq\gamma.$
B.3 Auxiliary Lemmas
Lemma B.6 (Oracle Switching Budget).
Under Assumption B.5, the number of switches of the context-indexed oracle action satisfies
@@ B.3 Auxiliary Lemmas
Here
$B_{T}:=\sum_{t=2}^{T}\mathbf{1}\{y_{t}^{*}(x_{t})\neq y_{t-1}^{*}(x_{t})\},$
where both oracle actions are evaluated at the same context $x_{t}$. This convention is used only for the switching analysis and is justified by the uniform contextual margin assumption.
@@
Fix $t\geq 2$ and abbreviate $y_{t}^{*}:=y_{t}^{*}(x_{t})$ and $y_{t-1}^{*}:=y_{t-1}^{*}(x_{t})$, the oracle-optimal action defined in B.2. If $y_{t}^{*}=y_{t-1}^{*}$, there is nothing to show. Suppose $y_{t}^{*}\neq y_{t-1}^{*}$. By optimality at time $t$,
@@
equivalently $u_{t-1}(x_{t},y_{t}^{*})-u_{t-1}(x_{t},y_{t-1}^{*})\leq-\gamma$. Thus, $\gamma\leq 2\phi_{\max}\|\theta_{t}-\theta_{t-1}\|_{2}$, which upon rearranging gives
@@
Lemma B.7 (Local variation bound).
@@
$\sum_{\tau\in W_{t}}\|\theta_{t}-\theta_{\tau}\|_{2}\leq WV_{t,W},\quad\sum^{T}_{t=1}V_{t,W}\leq WV_{T}.$
@@
Lemma B.8 (Self-normalized inequality).
Let
$\eta_{\tau}=p_{\tau}-\sigma(\langle\theta_{\tau},\psi_{\tau}\rangle)$
be conditionally zero-mean and 1-sub-Gaussian. Since $\|\psi_{\tau}\|_{2}\leq 2\phi_{\max}$, with probability at least $1-\delta$,
$\left\|\sum_{\tau\in W_{t}}\eta_{\tau}\psi_{\tau}\right\|_{2}\leq\sqrt{\lambda+4W\phi_{\max}^{2}}\sqrt{2\left[\frac{d}{2}\log\left(1+\frac{4W\phi_{\max}^{2}}{d\lambda}\right)+\log\left(\frac{1}{\delta}\right)\right]}.$
@@
We use the self-normalized vector martingale inequality following Abbasi-yadkori et al. [2011]. After rewriting the preference observation in pairwise form, the relevant feature vector is
$\psi_{\tau}=\phi(x_{\tau},y_{\tau}^{(1)})-\phi(x_{\tau},y_{\tau}^{(2)}),$
and the noise variable is
$\eta_{\tau}=p_{\tau}-\sigma(\langle\theta_{\tau},\psi_{\tau}\rangle),$
which is conditionally zero-mean under the Bradley–Terry preference model. The regularized design matrix is
$A_{t}=\sum_{\tau\in\mathcal{W}_{t}}\psi_{\tau}\psi_{\tau}^{\top}+\lambda I_{d}.$
Since $\|\phi(x,y)\|_{2}\leq\phi_{\max}$, we have
$\|\psi_{\tau}\|_{2}\leq\|\phi(x_{\tau},y_{\tau}^{(1)})\|_{2}+\|\phi(x_{\tau},y_{\tau}^{(2)})\|_{2}\leq 2\phi_{\max}.$
The self-normalized inequality gives, with probability at least $1-\delta$,
$\left\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\psi_{\tau}\right\|_{A_{t}^{-1}}\leq\sqrt{2\log\left(\frac{\det(A_{t})^{1/2}\det(\lambda I_{d})^{-1/2}}{\delta}\right)}.$
Now,
$\log\left(\frac{\det(A_{t})^{1/2}\det(\lambda I_{d})^{-1/2}}{\delta}\right)=\frac{1}{2}\log\left(\frac{\det(A_{t})}{\det(\lambda I_{d})}\right)+\log\left(\frac{1}{\delta}\right).$
Using the determinant-trace bound and $\|\psi_{\tau}\|_{2}\leq 2\phi_{\max}$,
$\det(A_{t})\leq\left(\lambda+\frac{1}{d}\sum_{\tau\in\mathcal{W}_{t}}\|\psi_{\tau}\|_{2}^{2}\right)^{d}\leq\left(\lambda+\frac{4W\phi_{\max}^{2}}{d}\right)^{d}.$
Since $\det(\lambda I_{d})=\lambda^{d}$, we obtain
$\frac{1}{2}\log\left(\frac{\det(A_{t})}{\det(\lambda I_{d})}\right)\leq\frac{d}{2}\log\left(1+\frac{4W\phi_{\max}^{2}}{d\lambda}\right).$
Therefore, with probability at least $1-\delta$,
$\left\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\psi_{\tau}\right\|_{A_{t}^{-1}}\leq\sqrt{2\left[\frac{d}{2}\log\left(1+\frac{4W\phi_{\max}^{2}}{d\lambda}\right)+\log\left(\frac{1}{\delta}\right)\right]}.$
Finally, converting from the $A_{t}^{-1}$ norm to the Euclidean norm gives
$\left\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\psi_{\tau}\right\|_{2}\leq\sqrt{\lambda_{\max}(A_{t})}\left\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\psi_{\tau}\right\|_{A_{t}^{-1}}.$
Using
$\lambda_{\max}(A_{t})\leq\lambda+\sum_{\tau\in\mathcal{W}_{t}}\|\psi_{\tau}\|_{2}^{2}\leq\lambda+4W\phi_{\max}^{2},$
we conclude that, with probability at least $1-\delta$,
$\left\|\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\psi_{\tau}\right\|_{2}\leq\sqrt{\lambda+4W\phi_{\max}^{2}}\sqrt{2\left[\frac{d}{2}\log\left(1+\frac{4W\phi_{\max}^{2}}{d\lambda}\right)+\log\left(\frac{1}{\delta}\right)\right]}.$
Lemma B.9 (Parameter estimation error bound).
Under Assumptions B.1 and B.2, let
$\eta_{\tau}=p_{\tau}-\sigma(\langle\theta_{\tau},\psi_{\tau}\rangle)$
be conditionally zero-mean and 1-sub-Gaussian. Then, for a fixed time $t$, with probability at least $1-\delta$,
$\|\widehat{\theta}_{t}-\theta_{t}\|_{2}\leq\frac{4L_{\sigma}\phi_{\max}^{2}}{m_{0}c}V_{t,W}+\frac{\sqrt{\lambda+4W\phi_{\max}^{2}}}{m_{0}cW}\sqrt{2\left[\frac{d}{2}\log\left(1+\frac{4W\phi_{\max}^{2}}{d\lambda}\right)+\log\left(\frac{1}{\delta}\right)\right]}+\frac{\lambda\theta_{\max}}{m_{0}cW}$
@@
Define the sliding-window logistic objective
$L_{t}(\theta)=\sum_{\tau\in\mathcal{W}_{t}}\left[\log(1+\exp(\theta^{\top}\psi_{\tau}))-p_{\tau}\theta^{\top}\psi_{\tau}\right]+\frac{\lambda}{2}\|\theta\|_{2}^{2}.$
Its gradient and Hessian are
$\nabla L_{t}(\theta)=\sum_{\tau\in\mathcal{W}_{t}}\left[\sigma(\theta^{\top}\psi_{\tau})-p_{\tau}\right]\psi_{\tau}+\lambda\theta,$
and
$\nabla^{2}L_{t}(\theta)=\sum_{\tau\in\mathcal{W}_{t}}\sigma(\theta^{\top}\psi_{\tau})\left(1-\sigma(\theta^{\top}\psi_{\tau})\right)\psi_{\tau}\psi_{\tau}^{\top}+\lambda I_{d}.$
By Assumptions B.1 and B.2,
$\lambda_{\min}(\nabla^{2}L_{t}(\theta))\geq m_{0}cW.$
@@
Using strong convexity and the optimality condition $\nabla L_{t}(\widehat{\theta}_{t})=0$,
$m_{0}cW\|\widehat{\theta}_{t}-\theta_{t}\|_{2}^{2}\leq\langle-\nabla L_{t}(\theta_{t}),\widehat{\theta}_{t}-\theta_{t}\rangle\leq\|\nabla L_{t}(\theta_{t})\|_{2}\|\widehat{\theta}_{t}-\theta_{t}\|_{2}.$
Thus,
$\|\widehat{\theta}_{t}-\theta_{t}\|_{2}\leq\frac{1}{m_{0}cW}\|\nabla L_{t}(\theta_{t})\|_{2}.$
Next, decompose the gradient at $\theta_{t}$:
$\nabla L_{t}(\theta_{t})=\sum_{\tau\in\mathcal{W}_{t}}\left[\sigma(\theta_{t}^{\top}\psi_{\tau})-\sigma(\theta_{\tau}^{\top}\psi_{\tau})\right]\psi_{\tau}-\sum_{\tau\in\mathcal{W}_{t}}\eta_{\tau}\psi_{\tau}+\lambda\theta_{t}.$
By the Lipschitz property of $\sigma$ and $\|\psi_{\tau}\|_{2}\leq 2\phi_{\max}$,
$\left\|\sum_{\tau\in\mathcal{W}_{t}}\left[\sigma(\theta_{t}^{\top}\psi_{\tau})-\sigma(\theta_{\tau}^{\top}\psi_{\tau})\right]\psi_{\tau}\right\|_{2}\leq\sum_{\tau\in\mathcal{W}_{t}}L_{\sigma}|\langle\theta_{t}-\theta_{\tau},\psi_{\tau}\rangle|\|\psi_{\tau}\|_{2}$
$\leq\sum_{\tau\in\mathcal{W}_{t}}L_{\sigma}\|\theta_{t}-\theta_{\tau}\|_{2}\|\psi_{\tau}\|_{2}^{2}\leq 4L_{\sigma}\phi_{\max}^{2}\sum_{\tau\in\mathcal{W}_{t}}\|\theta_{t}-\theta_{\tau}\|_{2}\leq 4L_{\sigma}\phi_{\max}^{2}WV_{t,W}.$
The noise term is bounded by Lemma B.8, and $\lambda\|\theta_{t}\|_{2}\leq\lambda\theta_{\max}$. Therefore,
$\|\nabla L_{t}(\theta_{t})\|_{2}\leq$
@@
$4L_{\sigma}\phi_{\max}^{2}WV_{t,W}+\sqrt{\lambda+4W\phi_{\max}^{2}}\sqrt{2\left[\frac{d}{2}\log\left(1+\frac{4W\phi_{\max}^{2}}{d\lambda}\right)+\log\left(\frac{1}{\delta}\right)\right]}+\lambda\theta_{\max}.$
Dividing by $m_{0}cW$ gives the result. ∎
Lemma B.10 (KL divergence bound).
Let $u_{t}(x_{t},y)=\langle\theta_{t},\phi(x_{t},y)\rangle$ and $\widehat{u}_{t}(x_{t},y)=\langle\widehat{\theta}_{t},\phi(x_{t},y)\rangle$. For the idealized KL-regularized policies induced by $u_{t}$ and $\widehat{u}_{t}$ with regularization coefficient $\beta_{\mathrm{ref}}$,
@@
$D_{\mathrm{KL}}\!\left(\pi^{\mathrm{kl}}_{t}(\cdot|x_{t})\,\|\,\pi_{t}(\cdot|x_{t})\right)\leq\frac{\phi_{\max}^{2}\|\theta_{t}-\widehat{\theta}_{t}\|_{2}^{2}}{2\beta_{\mathrm{ref}}^{2}}.$
@@
For the idealized KL-regularized comparator analyzed in this appendix, we use the exponential-tilted form
$\pi_{t}^{{kl}}(y_{t}|x_{t})=\frac{\pi_{\mathrm{ref},t}(y_{t}|x_{t})\exp(u_{t}(x_{t},y_{t})/\beta_{\mathrm{ref}})}{Z_{t}(x_{t})}.$
Similarly, the estimated policy induced by $\widehat{u}_{t}$ is
$\pi_{t}(y_{t}|x_{t})=\frac{\pi_{\mathrm{ref},t}(y_{t}|x_{t})\exp(\widehat{u}_{t}(x_{t},y_{t})/\beta_{\mathrm{ref}})}{\widehat{Z}_{t}(x_{t})},$
where $Z_{t}(x_{t})=\sum_{y_{t}}\pi_{ref,t}(y_{t}|x_{t})\exp(\frac{1}{\beta_{\mathrm{ref}}}u_{t}(x_{t},y_{t}))$ and $\widehat{Z}_{t}(x_{t})$ is defined analogously with $\widehat{u}_{t}$ in place of $u_{t}$. This idealized form is used only for the regret analysis and abstracts away from finite-sample sequence-level DPO optimization.
From the KL-divergence definition above, we first compute the log-ratio.
$\log{\frac{\pi^{kl}_{t}(y_{t}|x_{t})}{\pi_{t}(y_{t}|x_{t})}}=\log\frac{\frac{1}{Z(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta_{\mathrm{ref}}}u_{t}(x_{t},y_{t})}}{\frac{1}{\hat{Z}(x_{t})}\pi_{ref,t}(y_{t}|x_{t})\exp{\frac{1}{\beta_{\mathrm{ref}}}\hat{u}_{t}(x_{t},y_{t})}}$
$\log\frac{\hat{Z}(x_{t})}{Z(x_{t})}+\frac{1}{\beta_{\mathrm{ref}}}(u_{t}(x_{t},y_{t})-\hat{u}_{t}(x_{t},y_{t})).$